# Fluorescent Biomolecules Detectable in Near-Surface Ice on Europa

### v1

@@ 1.  Introduction
Recently, a novel probabilistic approach introduced the molecular assembly index (MA), a measure that quantifies the assembly complexity of molecules (Marshall et al., 2021). According to this measure, molecules with an MA score greater than 15 are extremely unlikely to form abiotically. Tyrosine and tryptophan have MA scores of 10 and 12, respectively, as opposed to glycine, for example, which has an MA score of 4. Thus, while the abiotic synthesis of the aromatic species is possible – it is not likely, which further elucidates the dearth of discovery of these species in abiotic settings. On Earth, the concentration of tryptophan has been measured to be about 0.1 part-per-billion (ppb) in the barren regions of the well-mixed ocean, where their synthesis is least likely, (Yamashita and Tanoue, 2004). The average abundances of tyrosine and phenylalanine are about 36 and 66 times higher, respectively (Moura et al., 2013). While the observed concentrations cannot be directly tied to specific synthesis mechanisms, we consider them a conservative estimate for the initial concentration of aromatic amino acids in a well-mixed Europan ocean that contains putative life.
@@ 2.1.  Radiolysis
We implemented a simplified approach to simulate the interaction of charged particles with near-surface ice, utilizing methods similar to those used in previous studies (Nordheim et al., 2018) (see Appendix A). Our model accounts for electrons and the three most abundant magnetospheric ions near Europa’s orbit: H+, O++, and S$+$ (Nordheim et al., 2022). Although there is some uncertainty regarding the charge states and their distribution among magnetospheric ions, our selection of these ions is primarily based on observational data (Mauk et al., 2004). To simplify the simulation process, we assumed a uniform bombardment pattern for the ions. While additional factors, such as an ocean-induced magnetic field (Kivelson et al., 2004), could be included, it has been shown that their impact on the bombardment pattern is relatively minor (Nordheim et al., 2022), while adding significant complexity to the simulation. Moreover, deviations from uniform ion bombardment are most pronounced in low-latitude regions where electron energy deposition dominates, and thus the effect of this non-uniformity is small when considering the total surface dose from all charged particles (Nordheim et al., 2022). The surface was modeled as pure water ice with a density of 0.5 g cm-3, incorporating the effects of porosity. This density is consistent with an expected porosity of approximately 0.3-0.5 for vapor-deposited water ice under Europan conditions (Mitchell et al., 2017).
@@
Electrons bombarding the trailing hemisphere are considerably more numerous but less energetic ($E<20$ MeV), resulting in the deposition rate of more than an order of magnitude more energy into the ice. Furthermore, these less energetic electrons exhibit smaller penetration depths, leading to increased radiolytic degradation of amino acids within the upper layers. Although energetic ions are sparser, they are also effective at degrading amino acids in the near-surface ice, because the two heavier ions deposit most of their energy within the top layer of the near-surface ice, in contrast to electrons, which tend to penetrate deeper (a detailed discussion can be found in Appendix A).
@@ 3.  Estimating Fluorescence Signal
Extinction of photon fluxes in vapor-deposited ice protects amino acids from photolytic degradation but also attenuates both the laser beam and the induced fluorescence traveling back to the surface (see Appendix E). Our findings indicate that such a laser can effectively penetrate the uppermost millimeter of vapor-deposited ice under Europan conditions, establishing this depth as the critical longevity depth scale of amino acids for significant fluorescence signal detection (see Appendix E). In comparison with reported detection thresholds of organic compounds buried in at least 2.4 cm of glacial ice in Earth-like conditions (Eshelman et al., 2019), Europa, and Enceladus to a lesser degree, present more challenging environments for applying such methodology, as the relevant layer prone for detection is also the one exposed to the harshest degradation.
To assess the detectability of aromatic amino acids through laser-induced spectroscopy, we assume uniformly distributed initial ppb-scale concentrations of tryptophan, tyrosine, and phenylalanine, with a mixing ratio of 1:36:66, respectively, within the ice. These concentrations represent conservative estimates derived from Earth’s oceans (Yamashita and Tanoue, 2004; Moura et al., 2013). Fig. 3 illustrates the net number of generated fluorescence photons, as a function of a 248.6 nm laser (Johnson and Hunter, 1980) energy and amino acid concentration, at three nominal terrain ages, demonstrating the rapid deterioration of the potential signal due to degradation, assuming a radiolytic constant of 0.034 MGy-1 (see Appendix E). The structure of fluorescence yield, which serves as a proxy for the degradation rate, is primarily controlled by the effectiveness of hemispheric radiolysis lenses, particularly in the trailing hemisphere. Secondary contributions arise from uniformly bombarding ions, with a lesser influence from photolysis, which is further modulated by the ice phase.
@@ 4.  Conclusion
We show that laser-induced fluorescence spectroscopy can effectively detect these biomolecules even from orbit. This technique can penetrate the uppermost millimeter of vapor-deposited ice, making it attractive for potential future missions targeting geologically young terrain, such as plume ejecta. Specifically, our findings underscore the importance of the polar regions on Europa as prime targets for biosignature detection, where preservation of aromatic amino acids may lead to detectable concentrations of these compounds. Future missions equipped with spectroscopic instruments will be well-positioned to explore these regions and potentially uncover evidence of biotic or abiotic synthesis of complex biomolecules. The ability to predict the presence of fluorescent organic molecules other than aromatic amino acids on Europa’s surface is critical for evaluating the feasbility of UV-induced fluorescence detection approach potential of aromatic amino acids. Detecting these coexisting molecules can, on the one hand, attenuate the unique fluorescence signatures of the aromatic amino acids and, on the other hand, provide additional context, revealing interactions between various organics and Europa’s radiation-altered ice. Such insights could help elucidate the history and origin of the extant organic material on Europa’s surface.
@@ Acknowledgements
The authors are grateful to the reviewers for their constructive reports, which helped improve the quality of the article.
@@ A.1.  Rationale
$k_{\rm{radio}}=\frac{0.05}{4.3}\cdot\frac{1}{103}+{4.3}\cdot\frac{36}{103}+0.05\cdot\frac{66}{103}+\frac{0.05}{\approx}0.034.$ (A.2)
We conducted particle transport simulations using G4beamline (Roberts and Kaplan, 2007), a GEANT4-based toolkit (Agostinelli et al., 2003), to model the depth-dependent energy deposition by energetic particles in water ice. The energy spectra for electrons and ions were derived from published fits (Mauk et al., 2004; Nordheim et al., 2018, 2022), based on Voyager and Galileo measurements (Paranicas et al., 2001). Depth profiles of deposited dose rates were computed for electrons, protons, and heavier ions (O++, S$+$) (see Figure A.1).
@@
Figure A.1: Integrated deposited dose rate from the bombardment of magnetospheric electrons and ions (H+, O++, and S$+$) for the leading (a) and trailing (b) hemispheres. The bombardment of ions is assumed to be uniform, and electron bombardment patterns follow a lens-like pattern centered on the equatorial region of each hemisphere.
@@ C.1.  Extinction Coefficients of Vapor-Deposited Ice
  • Strong Scattering Properties: The refractive indices and extinction coefficients of vapor-deposited water ice were recently measured (He et al., 2022) at different temperatures, distinguishing two regimes: amorphous ice at $T<130$ K and crystalline ice at $T>160$ K. The significant difference between the mean path length for UV-vis absorption (hundreds of meters) (Warren and Brandt, 2008) and scattering (tens of microns) (He et al., 2022) indicates that vapor-deposited ice can effectively attenuate incoming solar radiation, prolonging the lifetime of embedded molecules.
    Previous works demonstrate how optical properties of water ice change with different composition, deposition, and thermal evolution mechanisms (Warren and Brandt, 2008; Warren, 2019). For example, measured photolytic lifetimes of amino acids in crystalline ice indicated that the mean longevity of the considered species under Europan conditions is on the order of a decade (Orzechowska et al., 2007). In contrast, the optical properties of vapor-deposited ice, particularly its scattering coefficient, differ significantly from crystalline ice (He et al., 2022). Therefore, lifetime estimates for amino acids in Europan near-surface ice can vary considerably depending on the ice’s properties.
@@ Appendix D Amorphization and Crystallization
For crystallization, we solve the Avrami equation (with $n=2$ and $\Delta H=60\,\rm{kJ/mol}$ based on prior measurements (Jenniskens and Blake, 1996; Mitchell et al., 2017)), which describes the bulk of the crystallization process. The net crystallinity at any given location and time is derived by balancing the amorphization and crystallization rates.
@@ E.1.  Fluorescence Emission Attenuation and Potential Detection Strategies
Figure E.1: Degradation timescales of aromatic amino acids within vapor-deposited ice, allowing for an SNR of 3 detections of laser-induced fluorescence, for three radiolytic constants $k_{\rm{radio}}$: (a, b) 0.034 MGy-1 (c, d) 0.0034 MGy-1(e, f) 0.00034 MGy-1. Top: Detection from the surface. Bottom: Detection from orbit.

### v3

@@ 1.  Introduction
Recently, a novel probabilistic approach introduced the molecular assembly index (MA), a measure that quantifies the assembly complexity of molecules (Marshall et al., 2021). According to this measure, molecules with an MA score greater than 15 are extremely unlikely to form abiotically. Tyrosine and tryptophan have MA scores of 10 and 12, respectively, as opposed to glycine, for example, which has an MA score of 4. Thus, while the abiotic synthesis of the aromatic species is possible – it is not likely, which further elucidates the dearth of discovery of these species in abiotic settings. On Earth, the concentration of tryptophan has been measured to be about 0.1 parts-per-billion (ppb) in the barren regions of the well-mixed ocean, where its synthesis is least likely, (Yamashita and Tanoue, 2004). The average abundances of tyrosine and phenylalanine are about 36 and 66 times higher, respectively (Moura et al., 2013). While the observed concentrations cannot be directly tied to specific synthesis mechanisms, we consider them a conservative estimate for the initial concentration of aromatic amino acids in a well-mixed Europan ocean that contains putative life.
@@ 2.1.  Radiolysis
We implemented a simplified approach to simulate the interaction of charged particles with near-surface ice, utilizing methods similar to those used in previous studies (Nordheim et al., 2018) (see Appendix A). Our model accounts for electrons and the three most abundant magnetospheric ions near Europa’s orbit: H+, O2+, and S3+ (Nordheim et al., 2022). Although there is some uncertainty regarding the charge states and their distribution among magnetospheric ions, our selection of these ions is primarily based on observational data (Mauk et al., 2004). To simplify the simulation process, we assumed a uniform bombardment pattern for the ions. While additional factors, such as an ocean-induced magnetic field (Kivelson et al., 2004), could be included, it has been shown that their impact on the bombardment pattern is relatively minor (Nordheim et al., 2022), while adding significant complexity to the simulation. Moreover, deviations from uniform ion bombardment are most pronounced in low-latitude regions where electron energy deposition peaks, and thus the effect of this non-uniformity is small when considering the total surface dose from all charged particles (Nordheim et al., 2022). The surface was modeled as pure water ice with a density of 0.5 g cm-3, incorporating the effects of porosity. This density is consistent with an expected porosity of approximately 0.3-0.5 for vapor-deposited water ice under Europan conditions (Mitchell et al., 2017).
@@
Electrons bombarding the trailing hemisphere are considerably more numerous but less energetic ($\gtrsim 20$ MeV), resulting in the deposition rate of more than an order of magnitude more energy into the ice. Furthermore, these less energetic electrons exhibit smaller penetration depths, leading to increased radiolytic degradation of amino acids within the upper layers. Although energetic ions are sparser, they are also effective at degrading amino acids in the near-surface ice, because the two heavier ions deposit most of their energy within the top layer of the near-surface ice, in contrast to electrons, which tend to penetrate deeper (a detailed discussion can be found in Appendix A).
@@ 3.  Estimating Fluorescence Signal
Extinction of photon fluxes in vapor-deposited ice protects amino acids from photolytic degradation but also attenuates both the laser beam and the induced fluorescence traveling back to the surface (see Appendix E). Our findings indicate that such a laser can effectively penetrate the uppermost millimeter of vapor-deposited ice under Europan conditions, establishing this depth as the critical longevity depth scale of amino acids for significant fluorescence signal detection (see Appendix E). In comparison with reported detection thresholds of organic compounds buried in at least 2.4 cm of glacial ice in Earth-like conditions (Eshelman et al., 2019), Europa and Enceladus, to a lesser degree, present more challenging environments for applying such methodology, as the relevant layer prone for detection is also the one exposed to the harshest degradation.
To assess the detectability of aromatic amino acids through laser-induced spectroscopy, we assumed an initial concentration of tryptophan to be 0.1 ppb and a a mixing ratio of 1:36:66 for tryptophan, tyrosine, respectively. The three amino acids were modeled as uniformly-distributed across the vertical ice column. These concentrations represent conservative estimates derived from Earth’s oceans (Yamashita and Tanoue, 2004; Moura et al., 2013). Fig. 3 illustrates the net number of generated fluorescence photons, as a function of a 248.6 nm laser (Johnson and Hunter, 1980) energy and amino acid concentration, at three nominal terrain ages, demonstrating the rapid deterioration of the potential signal due to degradation, assuming a radiolytic constant of 0.034 MGy-1 (see Appendix E). The structure of fluorescence yield, which serves as a proxy for the degradation rate, is primarily controlled by the effectiveness of hemispheric radiolysis lenses, particularly in the trailing hemisphere. Secondary contributions arise from uniformly bombarding ions, with a lesser influence from photolysis, which is further modulated by the ice phase.
@@ 4.  Conclusion
We show that laser-induced fluorescence spectroscopy can effectively detect these biomolecules even from orbit. This technique can penetrate the uppermost millimeter of vapor-deposited ice, making it attractive for potential future missions targeting geologically young terrain, such as plume ejecta. Specifically, our findings underscore the importance of the polar regions on Europa as prime targets for biosignature detection, where preservation of aromatic amino acids may lead to detectable concentrations of these compounds. Future missions equipped with spectroscopic instruments will be well-positioned to explore these regions and potentially uncover evidence of biotic or abiotic synthesis of complex biomolecules. The ability to predict the presence of fluorescent organic molecules other than aromatic amino acids on Europa’s surface is critical for evaluating the feasibility of the UV-induced fluorescence detection approach for aromatic amino acids. Detecting these coexisting molecules can, on the one hand, attenuate the unique fluorescence signatures of the aromatic amino acids and, on the other hand, provide additional context, revealing interactions between various organics and Europa’s radiation-altered ice. Such insights could help elucidate the history and origin of the extant organic material on Europa’s surface.
@@ Acknowledgements
We are grateful to Erin Redwing, Prof. Imke de Pater, and Dr. Chris McKay for their early insights that helped shape the direction of this work. We are also grateful to the reviewers for their constructive reports, which helped improve the quality of the article.
@@ A.1.  Rationale
$k_{\rm{radio}}=\frac{0.05}{4.3}\cdot\frac{1}{103}+{4.3}\cdot\frac{36}{103}+0.05\cdot\frac{66}{103}+\frac{0.05}{4.3}\approx 0.034.$ (A.2)
We conducted particle transport simulations using G4beamline (Roberts and Kaplan, 2007), a GEANT4-based toolkit (Agostinelli et al., 2003), to model the depth-dependent energy deposition by energetic particles in water ice. The energy spectra for electrons and ions were derived from published fits (Mauk et al., 2004; Nordheim et al., 2018, 2022), based on Voyager and Galileo measurements (Paranicas et al., 2001). Depth profiles of deposited dose rates were computed for electrons, protons, and heavier ions (O2+, S3+) (see Figure A.1).
@@
Figure A.1: Integrated deposited dose rate from the bombardment of magnetospheric electrons and ions (H+, O2+, and S3+) for the leading (a) and trailing (b) hemispheres. The bombardment of ions is assumed to be uniform, and electron bombardment patterns follow a lens-like pattern centered on the equatorial region of each hemisphere.
@@ C.1.  Extinction Coefficients of Vapor-Deposited Ice
  • Strong Scattering Properties: The refractive indices and extinction coefficients of vapor-deposited water ice were recently measured at different temperatures (He et al., 2022), distinguishing two steady-state regimes: amorphous ice at $T\lesssim 130$ K and crystalline ice at $T\gtrsim 160$ K. The significant difference between the mean path length for UV-vis absorption (hundreds of meters) (Warren and Brandt, 2008) and scattering (tens of microns) (He et al., 2022) indicates that vapor-deposited ice can effectively attenuate incoming solar radiation, prolonging the lifetime of embedded molecules.
    Previous works demonstrated how optical properties of water ice change with different composition, deposition, and thermal evolution mechanisms (Warren and Brandt, 2008; Warren, 2019). For example, measured photolytic lifetimes of amino acids in crystalline ice indicated that the mean longevity of the considered species under Europan conditions is on the order of a decade (Orzechowska et al., 2007). In contrast, the optical properties of vapor-deposited ice, particularly its scattering coefficient, differ significantly from crystalline ice (He et al., 2022). Therefore, lifetime estimates for amino acids in Europan near-surface ice can vary considerably depending on the ice’s properties.
@@ Appendix D Amorphization and Crystallization
For crystallization, we solve the Avrami equation (with $n=2$ and $\Delta H=60\,\rm{kJ/mol}$ based on prior measurements (Jenniskens and Blake, 1996; Mitchell et al., 2017)), which describes the bulk of the crystallization process of porous ice. The net crystallinity at any given location and time is derived by balancing the amorphization and crystallization rates.
@@ E.1.  Fluorescence Emission Attenuation and Potential Detection Strategies
Figure E.1: Degradation timescales of aromatic amino acids within vapor-deposited ice, allowing for an SNR of 3 detection of laser-induced fluorescence, for three radiolytic constants $k_{\rm{radio}}$: (a, b) 0.034 MGy-1 (c, d) 0.0034 MGy-1(e, f) 0.00034 MGy-1. Top: Detection from the surface. Bottom: Detection from orbit.